Lindon Company is the exclusive distributor for an automotive product that sells for $54.00 per unit and has a CM ratio of 30%. The company’s fixed expenses are $388,800 per year. The company plans to sell 28,600 units this year. Required: 1. What are the variable expenses per unit? (Round your "per unit" answer to 2 decimal places.) 2. What is the break-even point in unit sales and in dollar sales? 3. What amount of unit sales and dollar sales is required to attain a target profit of $226,800 per year? 4. Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $5.40 per unit. What is the company’s new break-even point in unit sales and in dollar sales? What dollar sales is required to attain a target profit of $226,800?
Answer #1

CM ratio = Contribution Margin/Sales

Hence, contribution margin = 54*30% = $16.2

1.Variable cost per unit = Sales price per unit – contribution margin per unit

= 54-16.2

= $37.8

2.Break even point in unit sales = Total Fixed costs/Contribution Margin per unit

= 388,800/16.2

= 24,000 units

In Sales Dollars = Fixed costs/CM ratio

= 388,800/30%

= $1,296,000

3.Target Profit = $226,800

Add: Fixed costs = 388,800

Target Contribution Margin = $615,600

Units Required = 615,600/16.2

= 38,000 units

Sales dollars = 615,600/30%

= $2,052,000

4.Contribution margin = 54-32.4 = $21.6

Break even point = 388,800/21.6 = 18,000 units

And 18,000*54 = $972,000

Target Profit = 226,800

Desired contribution margin = $615,600

Dollar Sales = 615,600/40%

= $1,539,000